OpenStudy (anonymous):

How would you solve 4y-x=8 graphically and what is slope and y intercept

Mehek (mehek14):

rewrite to the general equation y=mx+b whereas m = slope and b=y-intercept 4y-x=8 4y = x + 8 y = 4x + 2 m = 4 b = 2 you can solve this graphically by substituting the x and y values by the points

OpenStudy (anonymous):

Ight thanks dude